Let X equal to the number of heads after 4 flips of a fair coin? Derive the probability mass function for X, and plot it. Also, compute the E[X] of X.
Answer #1

The probability of 0 heads is equal to the probability of 4 heads.

Since the coin is fair, the probability of no heads in 4 tosses is

rac{1}{2} imes rac{1}{2} imes rac{1}{2} imes rac{1}{2} = rac{1}{16}

Same is the probability of getting 4 heads.

The probability of getting 1, 2, or 3 heads is different. For example, there are four ways to get 1 head:

HTTT, THTT, TTHT, TTTH

Each of these cases has probability 1/16 as before, but there are 4 of them so the total probability of getting 1 head is 4/16.

It is a similar procedure to count the number of ways to get 2 or 3 heads. More generally, the number of ways to get x heads out of n coin tosses is given by the binomial coefficient:

nCr (n, r)

To generalize even further, this coin tossing experiment follows the binomial distribution. If the probability of tossing a head is p then the PMF is given by

P(X = x) = {n choose x} p^{x}(1-p)^{n-x}

Since p=1/2

We have

P(X-) 0.51 0.5)

P(X 0) -Ί 0,5°(1- 0,5 0 ) )4-0-1/16

P(X = 1) = {4 choose 1} 0.5^{1}(1-0.5)^{4-1}=4/16

P(X = 2) = {4 choose 2} 0.5^{2}(1-0.5)^{4-2}=6/16

P(X = 3) = {4 choose 3} 0.5^{3}(1-0.5)^{4-3}=4/16

P(X = 4) = {4 choose 4} 0.5^{4}(1-0.5)^{4-4}=1/16

Expectation of X

P(x)x*P(x) O.iNi 1.22 0.38 1.22 O.iNi Sum 1.22 0.75 0.75 0.25

E(X)=Sum(x*P(x))=2